# On-the-Fly OVD Adaptation with FLAME: Few-shot Localization via Active Marginal-Samples Exploration

### v1

@@ Abstract
Open-vocabulary object detection (OVD) models offer remarkable flexibility by detecting objects from arbitrary text queries. However, their zero-shot performance in specialized domains like Remote Sensing (RS) is often compromised by the inherent ambiguity of natural language, limiting critical downstream applications. For instance, an OVD model may struggle to distinguish between fine-grained classes such as "fishing boat" and "yacht" since their embeddings are similar and often inseparable. This can hamper specific user goals, such as monitoring illegal fishing, by producing irrelevant detections. To address this, we propose a cascaded approach that couples the broad generalization of a large pre-trained OVD model with a lightweight few-shot classifier. Our method first employs the zero-shot model to generate high-recall object proposals. These proposals are then refined for high precision by a compact classifier trained in real-time on only a handful of user-annotated examples - drastically reducing the high costs of RS imagery annotation. The core of our framework is FLAME, a one-step active learning strategy that selects the most informative samples for training. FLAME identifies, on the fly, uncertain marginal candidates near the decision boundary using density estimation, followed by clustering to ensure sample diversity. This efficient sampling technique achieves high accuracy without costly full-model fine-tuning and enables instant adaptation, within less then a minute, which is significantly faster than state-of-the-art alternatives. Our method consistently surpasses state-of-the-art performance on RS benchmarks, establishing a practical and resource-efficient framework for adapting foundation models to specific user needs.
@@ 1 Introduction
The recent advancements in large-scale vision-language models (VLMs) such as CLIP [21] have catalyzed a paradigm shift in computer vision, giving rise to Open-Vocabulary Object Detection (OVD) [29]. Unlike traditional detectors limited to predefined categories, OVD models can identify objects described by arbitrary natural language text, offering unprecedented flexibility. This is particularly transformative for remote sensing (RS), where cataloging every possible class is intractable. Early OVD methods adapted standard detectors by replacing the classifier head with text embeddings [11], leveraging the semantic richness of VLMs to generalize to unseen categories. However, the inherent ambiguity of text queries often leads to significant drops in precision, limiting the utility of pure zero-shot systems.
Figure 1: A visual demonstration of performance improvement from Zero-Shot to Few-Shot detection using DIOR dataset [30]. The Zero-Shot model (center) produces noisy and unreliable results, identifying the ’chimneys’ but with low confidence and accompanied by several false positives. Our Few-Shot method (right) refines this output, successfully eliminating the false positives and accurately detecting all four chimneys shown in the Ground Truth (left).
To overcome the limitations of pure zero-shot systems, one alternative is Few-Shot Object Detection (FSOD) [14], which adapts models to novel categories using only a handful of annotated examples. In RS, FSOD is critical due to the difficulty and cost of acquiring dense labels[1]. While effective, common FSOD strategies like meta-learning or fine-tuning [26] can be computationally intensive. To address this, Parameter-Efficient Fine-Tuning (PEFT) techniques such as LoRA [12] have emerged to alleviate these costs by reducing the number of trainable parameters.
@@
A distinct paradigm explores a hybrid approach that merges OVD and FSOD, using few-shot supervision to enhance and expand an open-vocabulary detector’s existing knowledge within a single, unified framework [6]. Several strategies explore this hybrid model: prompt-based methods [10, 31] learn continuous prompts from support sets to improve category alignment, while Transformer-based methods like OV-DETR [28] and OWL-ViT [19] show strong generalization.
The success of these hybrid approaches, which use only a handful of examples, hinges on the efficient selection of the most informative ones. This challenge is addressed by Active Learning (AL) [24], which queries an oracle for the most beneficial labels. Common AL strategies include uncertainty-based sampling [16], diversity-based sampling [23], or their combinations [7]. Building on this foundation, our work proposes a cascaded OVD–FSOD framework with a novel AL strategy specifically designed to resolve semantic ambiguity in RS imagery efficiently and effectively.
@@ 2 Method
Theory and Motivations.
Our framework lays on the observation that a binary classifier, whether an SVM [25] or a positively homogeneous neural network [20], can be determined entirely by its margin (support) examples. Equivalently, if one removes all non-support training points and retrains, the resulting classifier is unchanged. Building on this, our few-shot procedure identifies a small set of near-boundary examples (the "few-shots"), asks the user to label them, and trains a lightweight model on the fly. Despite using only a handful of points, this model matches the classifier that would have been obtained from training on the full dataset, which may be too large or impractical for real-time training. The Lemma below formalizes this fact for the SVM. Similarly, we prove the same for soft margin (SVM) and for neural networks in Appendix B. The proofs of the Lemmas are also relegated to the Appendix B.
Lemma 2.1 (Support–determination for hard–margin SVM).
$\min_{w,b}\ \tfrac{1}{2}\|w\|^{2}\quad\text{s.t.}\quad y_{i}\,(w^{\top}x_{i}+b)\ \geq\ 1\quad(i=1,\dots,n).$ (P)
  1. 1.
    $\min_{w,b}\ \tfrac{1}{2}\|w\|^{2}\quad\text{s.t.}\quad y_{i}\,(w^{\top}x_{i}+b)\ \geq\ 1\quad(i\in S).$ ($\mathrm{P}_{S}$)
  2. 2.
[…]
Remark 2.2 (Kernel SVM).
Marginal Samples Retrieval.
We propose a one-stage active learning strategy that pinpoints the most informative samples for training a lightweight, class-specific binary classifier. This algorithm 1 allows a large-scale, zero-shot OVD model to be adapted to a new target class efficiently, in real-time, and with minimal human supervision. The method is illustrated in Figure 2.
@@
First, we identify uncertain candidates by augmenting image embeddings with their zero-shot similarity to the text query and applying density estimation in a projected (PCA) augmented-embedding-space. Samples at the distribution’s margins are retained as they carry the most informative ambiguity. From this pool, we promote diversity by clustering and selecting one representative per cluster, yielding $K$ candidate shots for annotation. The user then labels these few informative samples, forming an initial dataset. To mitigate imbalance, we apply Synthetic Minority Over-sampling Technique (SMOTE) [4] for extremely fast augmentation. This procedure would contribute to a balanced, representative, and efficiently training to take place shortly after.
Finally, using the (augmented) few-shots returned by Algorithm 1, we train a compact classifier, by default an Radial Basis kernel (RBF) SVM [22], which is trained to find a non-linear separating hyperplane. Note that our efficient framework could support many lightweight alternatives such as: Two-Layer Multi-Layer Perceptron (MLP) under binary cross-entropy loss function, or encoder-classifier with Triplet Loss [9].
3 Experiments
To evaluate its performance, our few-shot method is benchmarked against a zero-shot baseline and leading state-of-the-art approaches, as summarized in Table 1. To that end, we leverage the following two RS datasets: (1) DOTA [27] (Dataset for Object Detection in Aerial Images): A large-scale RS dataset with multi-class, multi-oriented objects annotated in high-resolution aerial images for object detection. (2) DIOR[17] (Dataset for Object Detection in Optical RS Images): A diverse large-scale dataset of optical RS images containing numerous object categories across varying conditions and resolutions for robust detection.
Method DOTA DIOR
Zero-shot OWL-ViT-v2 (Baseline) 13.774% 14.982%
Zero-shot RS-OWL-ViT-v2 31.827% 29.387%
Jeune et. al [15] 37.1% 35.6%
SIoU [13] 45.88% 52.85%
Prototype-based FSOD with DINOv2 [3] 41.40% 26.46%
FLAME cascaded on RS-OWL-ViT-v2 53.96% 53.21%
We first evaluate the zero-shot performance of the baseline OWL-ViT-v2 model [18], which was pre-trained on the vast, generic multilingual WebLI dataset [5]. We then consider the RS-OWL-ViT-v2 model [2], a remote sensing variant of OWL-ViT-v2 fine-tuned on the RS-WebLI dataset [2], which consists of three million aerial and satellite images from the original WebLI dataset and on a collection of $67,000$ aerial images annotated for remote sensing object detection across $34$ categories. This improved zero-shot performance model serves as the starting point for FLAME.
In addition to demonstrating the highest performance, our method adapts in approximately $1$ minute for each label on a standard CPU. This is a considerable advantage over fine-tuning approaches that typically require GPU machines and several hours to complete. Further experiments can be found in section Appendix C. Illustration of the algorithm is presented in Figure 1.
4 Discussion
This work presents a practical and resource-efficient paradigm for adapting large foundational models to specific user needs in real time. By confining the adaptation to a lightweight classifier trained on actively selected samples, our method avoids costly fine-tuning while effectively resolving semantic ambiguities. This on-the-fly approach paves the way for more dynamic, user-in-the-loop applications in RS, OVD, and other specialized domains.
Appendix A Illustrations
Figure 2: Overview of the proposed few-shot sampling method. The method is followed by the stages: (1) uncertainty-based filtering using density estimation to identify ambiguous candidates near the decision boundary, (2) clustering-based diversity sampling to ensure representative coverage, (3) interactive user annotation of the selected samples, (4) conditional data augmentation with SMOTE or SVM-SMOTE to balance classes, and (5) lightweight classifier training (e.g., SVM or MLP) on the augmented set. This cascaded process refines the zero-shot proposals from a large open-vocabulary detector into an accurate, real-time few-shot classifier without full-model fine-tuning
Appendix B Proofs
Proof of Lemma 2.1.
Introduce Lagrange multipliers $\alpha_{i}\geq 0$ for the constraints in (P). The Lagrangian is
$\mathcal{L}(w,b,\alpha)=\tfrac{1}{2}\|w\|^{2}-\sum_{i=1}^{n}\alpha_{i}\bigl(y_{i}(w^{\top}x_{i}+b)-1\bigr).$
The KKT conditions for (P) are:
(stationarity) $\nabla_{w}\mathcal{L}=w-\sum_{i=1}^{n}\alpha_{i}y_{i}x_{i}=0,$ (1)
$\nabla_{b}\mathcal{L}=-\sum_{i=1}^{n}\alpha_{i}y_{i}=0,$ (2)
(primal feasibility) $y_{i}(w^{\top}x_{i}+b)\geq 1\quad(\forall i),$ (3)
(dual feasibility) $\alpha_{i}\geq 0\quad(\forall i),$ (4)
(complementary slackness) $\alpha_{i}\bigl(y_{i}(w^{\top}x_{i}+b)-1\bigr)=0\quad(\forall i).$ (5)
Let $(w^{\star},b^{\star},\alpha^{\star})$ be a KKT triple for (P). By complementary slackness (5), if $i\notin S$ then $y_{i}(w^{\star\top}x_{i}+b^{\star})>1$ and hence $\alpha_{i}^{\star}=0$. Therefore (1)–(2) reduce to
and feasibility plus (5) hold with indices restricted to $S$. These are precisely the KKT conditions for the reduced problem ($P_{S}$). This proves item (1).
@@ Appendix B Proofs
Combining (1) and (2), the optimal solution set of (P) coincides with that of ($P_{S}$). Since the primal objective $\tfrac{1}{2}\|w\|^{2}$ is strictly convex in $w$, the optimal $w^{\star}$ is unique; hence the decision function $x\mapsto\operatorname{sign}(w^{\star\top}x+b^{\star})$ is unchanged when the non–support training points are removed. $\Box$
Lemma B.1 (Support examples–determination for homogeneous networks).
@@
Proof of Lemma B.1.
The KKT conditions for (6) at $(\theta,\lambda)$ are:
stationarity: $\nabla_{\theta}\!\left(\tfrac{1}{2}\|\theta\|^{2}-\sum_{i=1}^{n}\lambda_{i}\,y_{i}\,\Phi(\theta;x_{i})\right)=0,$ (8)
primal feasibility: $y_{i}\,\Phi(\theta;x_{i})\geq 1\quad(\forall i),$ (9)
dual feasibility: $\lambda_{i}\geq 0\quad(\forall i),$ (10)
complementary slackness: $\lambda_{i}\big(y_{i}\,\Phi(\theta;x_{i})-1\big)=0\quad(\forall i).$ (11)

### v2

@@ Abstract
Open-vocabulary object detection (OVD) models offer remarkable flexibility by detecting objects from arbitrary text queries. However, their zero-shot performance in specialized domains like Remote Sensing is often compromised by the inherent ambiguity of natural language, limiting critical downstream applications. For instance, an OVD model may struggle to distinguish between fine-grained classes such as "fishing boat" and "yacht" since their embeddings are similar and often hard to separate. This can hamper specific user goals, such as monitoring illegal fishing, by producing irrelevant detections. To address this, we propose a cascaded approach that couples the broad generalization of a large pre-trained OVD model with a lightweight few-shot classifier. Our method first employs the zero-shot model to generate high-recall object proposals. These proposals are then refined for high precision by a compact classifier trained in real-time on only a handful of user-annotated examples - drastically reducing the high costs of remote sensing imagery annotation. The core of our framework is FLAME, a one-step active learning strategy that selects the most informative samples for training. FLAME identifies, on the fly, uncertain marginal candidates near the decision boundary using density estimation, followed by clustering to ensure sample diversity. This efficient sampling technique achieves high accuracy without costly full-model fine-tuning and enables instant adaptation, within less than a minute, which is significantly faster than state-of-the-art alternatives. Our method consistently surpasses state-of-the-art performance on remote sensing benchmarks, establishing a practical and resource-efficient framework for adapting foundation models to specific user needs.
Figure 1: A visual demonstration of performance improvement from Zero-Shot to Few-Shot detection using DIOR dataset [28]. The Zero-Shot model (center) produces noisy and unreliable results, identifying the ’chimneys’ but with low confidence and accompanied by several false positives. Our Few-Shot method (right) refines this output, successfully eliminating the false positives and accurately detecting all four chimneys shown in the Ground Truth (left). Similarly, the bottom row showcases the detection of a ’dam’. The Zero-Shot model struggles with false positives such as ’bridge’ and ’harbor’, which are corrected by the more precise Few-Shot approach.
@@ 1 Introduction
The recent advancements in large-scale vision-language models (VLMs) such as CLIP [19] have catalyzed a paradigm shift in computer vision, giving rise to Open-Vocabulary Object Detection (OVD) [27]. Unlike traditional detectors limited to predefined categories, OVD models can identify objects described by arbitrary natural language text, offering unprecedented flexibility. This is particularly transformative for remote sensing, where cataloging every possible class is intractable. Early OVD methods adapted standard detectors by replacing the classifier head with text embeddings [10], leveraging the semantic richness of VLMs to generalize to unseen categories. However, the inherent ambiguity of text queries often leads to significant drops in precision, limiting the utility of pure zero-shot systems.
To overcome the limitations of pure zero-shot systems, one alternative is Few-Shot Object Detection (FSOD) [13], which adapts models to novel categories using only a handful of annotated examples. In remote sensing, FSOD is critical due to the difficulty and cost of acquiring dense labels[1]. While effective, common FSOD strategies like meta-learning or fine-tuning [24] can be computationally intensive. To address this, Parameter-Efficient Fine-Tuning (PEFT) techniques such as LoRA [11] have emerged to alleviate these costs by reducing the number of trainable parameters.
@@
A distinct paradigm explores a hybrid approach that merges OVD and FSOD, using few-shot supervision to enhance and expand an open-vocabulary detector’s existing knowledge within a single, unified framework [5]. Several strategies explore this hybrid model: prompt-based methods [9, 29] learn continuous prompts from support sets to improve category alignment, while Transformer-based method like OV-DETR [26] shows strong generalization.
The success of these hybrid approaches, which use only a handful of examples, hinges on the efficient selection of the most informative ones. This challenge is addressed by Active Learning (AL) [22], which queries an oracle for the most beneficial labels. Common AL strategies include uncertainty-based sampling [15], diversity-based sampling [21], or their combinations [6]. Building on this foundation, our work proposes a cascaded OVD–FSOD framework with a novel AL strategy specifically designed to resolve semantic ambiguity in remote sensing imagery efficiently and effectively.
Figure 2: Overview of the proposed few-shot sampling method. The method follows the stages: (1) uncertainty-based filtering using density estimation to identify ambiguous candidates near the decision boundary, (2) clustering-based diversity sampling to ensure representative coverage, (3) interactive user annotation of the selected samples, (4) conditional data augmentation with SMOTE or SVM-SMOTE to balance classes, and (5) lightweight classifier training (e.g., SVM or MLP) on the augmented set. This cascaded process refines the zero-shot proposals from a large open-vocabulary detector into an accurate, real-time few-shot classifier without full-model fine-tuning
@@ 2 Method
Theory and Motivations.
Our framework lays on the observation that a binary classifier, whether an SVM [23] or a positively homogeneous neural network [18], can be determined entirely by its margin (support) examples. Equivalently, if one removes all non-support training points and retrains, the resulting classifier is unchanged. Building on this, our few-shot procedure identifies a small set of near-boundary examples (the "few-shots"), asks the user to label them, and trains a lightweight model on the fly. Despite using only a handful of points, this model matches the classifier that would have been obtained from training on the full dataset, which may be too large or impractical for real-time training. The lemmas below formalize this fact for the SVM, soft margin SVM and for neural networks. The proofs of the Lemmas are relegated to the Appendix A.
Lemma 2.1 (Support–determination for hard–margin SVM).
$\min_{w,b}\ \tfrac{1}{2}\|w\|^{2}\;\text{s.t.}\;y_{i}\,(w^{\top}x_{i}+b)\ \geq\ 1,\;(i=1,\dots,n).$ (P)
  1. 1.
    $\min_{w,b}\ \tfrac{1}{2}\|w\|^{2}\quad\text{s.t.}\;y_{i}\,(w^{\top}x_{i}+b)\ \geq\ 1,\;(i\in S).$ ($\mathrm{P}_{S}$)
  2. 2.
@@
Remark 2.2 (Kernel SVM).
Our claim for the non-separable embeddings case, which is the soft marginal SVM, is stated in the following lemma.
Lemma 2.2 (Support-determination for soft-margin SVM).
$\min_{w,b,\xi}$ $\quad\frac{1}{2}\|w\|^{2}+C\sum_{i=1}^{n}\xi_{i}$
$\text{s.t.}\quad y_{i}(w^{\top}x_{i}+b)\geq 1-\xi_{i},\quad(i=1,\dots,n)$ ($P$)
  1. 1.
    $\min_{w,b,\xi}\quad\frac{1}{2}\|w\|^{2}+C\sum_{i=1}^{n}\xi_{i}$
    $\quad\text{s.t.}\quad y_{i}(w^{\top}x_{i}+b)\geq 1-\xi_{i},\quad(i\in S)$ ($P_{S}$)
  2. 2.
[…]
Consequently, (2.2) and (1) have the same optimal solutions $(w,b)$. Retraining the soft-margin SVM after removing all non-support points ($i\notin S$) leaves the classifier $x\mapsto\operatorname{sign}(w^{\top}x+b)$ unchanged.
Lastly, the following Lemma formalizes our claim for the case of neural network.
Lemma 2.3 (Support examples–determination for homogeneous networks).
@@
Method DOTA DIOR
Zero-shot OWL-ViT-v2 (Baseline) 13.774% 14.982%
Zero-shot RS-OWL-ViT-v2 31.827% 29.387%
Jeune et. al [14] 37.1% 35.6%
SIoU [12] 45.88% 52.85%
Prototype-based FSOD with DINOv2 [2] 41.40% 26.46%
FLAME cascaded on RS-OWL-ViT-v2 53.96% 53.21%
Marginal Samples Retrieval.
We propose a one-stage active learning strategy that pinpoints the most informative samples for training a lightweight, class-specific binary classifier. Algorithm 1 allows a large-scale, zero-shot OVD model to be adapted to a new target class efficiently, in real-time, and with minimal human supervision. The method is illustrated in Figure 2. First, we identify uncertain candidates by augmenting image embeddings with their zero-shot similarity to the text query and applying density estimation in a projected (PCA) augmented-embedding-space. Samples at the distribution’s margins are retained as they carry the most informative ambiguity. From this pool, we promote diversity by clustering and selecting one representative per cluster, yielding $K$ candidate shots for annotation. The user then labels these few informative samples, forming an initial dataset. To mitigate imbalance, we apply Synthetic Minority Over-sampling Technique (SMOTE) [3] for extremely fast augmentation. This procedure would contribute to a balanced, representative, and efficient training to take place shortly after.
DIOR Dataset Class Zero Shot Few Shot expressway service area 0.03 0.82 expressway toll station 0 0.99 airplane 0.84 0.99 airport 0 baseball field 0.62 0.93 basketball court 0.66 0.87 bridge 0.21 0.49 chimney 0.11 0.94 dam 0.04 0.71 golf field 0.01 0.72 ground track field 0.5 0.79 harbor 0.33 0.64 overpass 0.1 0.75 ship 0.72 0.93 stadium 0.57 0.86 storage tank 0.73 0.68 tennis court 0.8 0.57 train station 0.01 vehicle 0.25 0.79 windmill 0.67 1
DOTA Dataset Class Zero Shot Few Shot Baseball Diamond 0.32 0.88 Basketball Court 0.56 0.83 Bridge 0.09 0.28 Container Crane 0.03 0.95 Ground Track Field 0.4 0.68 Harbor 0.36 0.82 Helicopter 0.39 0.73 Large Vehicle 0.32 0.87 Plane 0.78 0.54 Roundabout 0.24 0.91 Ship 0.71 0.82 Small Vehicle 0.28 0.77 Soccer Ball Field 0.48 0.77 Storage Tank 0.79 0.55 Swimming Pool 0.71 0.58 Tennis Court 0.77 0.01
Finally, using the (augmented) few-shots returned by Algorithm 1, we train a compact classifier, by default a Radial Basis kernel (RBF) SVM [20], which is trained to find a non-linear separating hyperplane. Note that our efficient framework could support many lightweight alternatives such as: Two-Layer Multi-Layer Perceptron (MLP) under binary cross-entropy loss function, or encoder-classifier with Triplet Loss [8]. Illustration schema of the algorithm is presented in Figure 2.
3 Experiments
To evaluate its performance, our few-shot method is benchmarked against a zero-shot baseline and leading state-of-the-art approaches, as summarized in Table 1. To that end, we leverage the following two remote sensing datasets: (1) DOTA [25] (Dataset for Object Detection in Aerial Images): A large-scale remote sensing dataset with multi-class, multi-oriented objects annotated in high-resolution aerial images for object detection; (2) DIOR [16] (Dataset for Object Detection in Optical remote sensing Images): A diverse large-scale dataset of optical remote sensing images containing numerous object categories across varying conditions and resolutions for robust detection.
We first evaluate the zero-shot performance of the baseline OWL-ViT-v2 model [17], which was pre-trained on the vast, generic multilingual WebLI dataset [4]. We then consider the RS-OWL-ViT-v2 model, a remote sensing variant of OWL-ViT-v2 fine-tuned on the RS-WebLI dataset [1], which consists of three million aerial and satellite images from the original WebLI dataset and on a collection of $67,000$ aerial images annotated for remote sensing object detection across $34$ categories. This improved zero-shot performance model serves as the starting point for FLAME.
Table 1 demonstrates that the FLAME cascaded on RS-OWL-VIT-v2 method achieves the highest Average Precision (AP) on both the DOTA ($53.96\%$) and DIOR ($53.21\%$) datasets among all compared Few-Shot Object Detection (FSOD) models.This superior performance is coupled with a significantly faster adaptation time (approximately 1 minute per label on a CPU) compared to competing fine-tuning approaches that typically require a hardware accelerator (such as TPU or GPU) and several hours.
Following, Table 2 provides a detailed per-class breakdown of the Average Precision (AP) on both the DIOR and DOTA datasets, comparing our few-shot method against the zero-shot baseline using the Zero-shot RS-OWL-VIT-v2 fine-tuned on RS-WebLI (which appear in second line of Table 1). The missing values in the ’Few-Shot’ columns indicate instances where the initial zero-shot retrieval step failed to find any relevant image embeddings. Without these initial candidates, the few-shot selection process could not proceed, resulting in a method failure for those specific classes. The Table demonstrates the substantial performance gains achieved by the Few-Shot (FLAME) method over the Zero-Shot baseline across a wide range of object categories on both the DIOR and DOTA datasets. For instance, the Few-Shot method dramatically improves AP for challenging classes like ’expressway toll station’ on DIOR (from $0\%$ to $99\%$) and ’Container Crane’ on DOTA (from $3\%$ to $95\%$), showcasing its effectiveness in resolving semantic ambiguity
4 Discussion
Remote sensing is a field that involves the acquisition of information about an object or area without making physical contact with it, typically using sensors on platforms such as satellites or aircraft. The proposed method provides a practical and resource-efficient framework for adapting foundational remote sensing OVD models to specific user needs. The cascaded architecture combines a large, pre-trained OVD model with a lightweight, few-shot classifier. This approach generates initial object-embedding proposals using the frozen weights of the zero-shot model, which are then refined by a compact classifier trained in real-time on a handful of user-annotated examples. This process drastically reduces annotation overhead while achieving high accuracy without the costly process of full-model fine-tuning. The core contribution is an efficient one-step active learning strategy that selects the most informative samples for user annotation. This strategy identifies a small number of uncertain candidates near the decision boundary using density estimation and then applies clustering to ensure a diverse training set. The method is designed to address the semantic ambiguity of text queries that hampers the zero-shot performance of pre-trained models.
Appendix A Proofs
Proof of Lemma 2.1.
Introduce multipliers $\alpha_{i}\geq 0$ for the constraints in (P). The Lagrangian is
$\mathcal{L}(w,b,\alpha)=\tfrac{1}{2}\|w\|^{2}-\sum_{i=1}^{n}\alpha_{i}\bigl(y_{i}(w^{\top}x_{i}+b)-1\bigr),$
and the KKT conditions are
(stationarity) $w=\sum_{i=1}^{n}\alpha_{i}y_{i}x_{i},\qquad\sum_{i=1}^{n}\alpha_{i}y_{i}=0,$
(primal feas.) $y_{i}(w^{\top}x_{i}+b)\geq 1\quad(\forall i),$
(dual feas.) $\alpha_{i}\geq 0\quad(\forall i),$
(comp. slackness) $\alpha_{i}\bigl(y_{i}(w^{\top}x_{i}+b)-1\bigr)=0\quad(\forall i).$